Write a quadratic equation in standard form with the solution set {2-sqrt5, 2+sqrt 5}.
The quadratic equation in standard form is ___.
(Simplify your answer. Type an equation using x as the variable.)

Answer by stanbon(75887)

Write a quadratic equation in standard form with the solution set
{2-sqrt5, 2+sqrt 5}.
----
f(x) = [x-(2-sqrt(t)][x-(2+sqrt(5)]
------
Rearrange:
f(x) = [(x-2)+sqrt(5)][(x-2)-sqrt(5)]
---------
Multiply:
------
f(x) = (x-2)^2 - 5
----
f(x) = x^2 -4x +4 -5
----
f(x) = x^2 -4x -1
========================
Cheers,
Stan H.
=====================